OpenStudy (anonymous):

Solve 2x2 – 5x + 1 = 3 for x.

OpenStudy (anonymous):

It is not clear Is that \[2x ^{2}\]

OpenStudy (solomonzelman):

\[2x^2 – 5x + 1 = 3\]\[2x^2 – 5x + 1 -3= 3-3\]\[2x^2 – 5x -2= 0\]then use the quadratic formula.